Find an equation of the line that passes through the given point and has the indicated slope m.
(8, 2), m = 1/4

show the graph.

Answer by Alan3354(69443)

Find an equation of the line that passes through the given point and has the indicated slope m.
(8, 2), m = 1/4
-------------
y - y1 = m*(x - x1) where
m is the slope
(x1,y1) is the point